omg inc has 5 million shares of common stock outstanding, 4 million shares of preferred stock outstanding and 6000 bonds. suppose the common share are selling for 25 per share, the preferred shares are selling for 24 per share, and the bonds are selling for 109 percent of par. what weight should use for debt in the computation of omg wacc
Answer #1

Total value of:

Common stock=(5*25)=$125 million

Preferred stock=(4*24)=$96 million

Bonds=6000*(1000*109%)=$6.54 million

Total value=(125+96+6.54)=$227.54 million

Hence weight of debt=(6.54/227.54)

=2.87%(Approx).